bill533:

The equation of line k is y = −1/5x − 1/8. What is the equation of the line that is perpendicular to line k and passes through (0, 0)? A. y = 5x B. x = -5x C. x = 8x D. y = -8x

bill533:

@563blackghost

563blackghost:

Perpendicular lines have reciprocal slopes (so you would flip the fraction and the sign). \(\large\bf{-\frac{1}{5}~~~\rightarrow~~~-\frac{\color{red}{5}}{\color{red}{1}}~~~\rightarrow~~~\frac{5}{1}}\) So your equation is: \(\large\bf{y=5x}\). Since your point is `(0,0)` then `b = 0`.

bill533:

A.

bill533:

Monica graphed the line y = 3 x + 2. Janice graphs a second line that is parallel to Monica’s and passes through (-2, 0). What is the equation of Janice’s line? What is the error in Rita’s reasoning? A. y = 3x + 6 B. y = −1/3x + 2 C. y = −1/3x − 2/3 D. y = 3x - 6

563blackghost:

Parallel lines have like slopes so your equation is: \(\bf{y=3x+b}\) Plug in your known point to find b..

bill533:

b = -3x + y

563blackghost:

you need to plug in your point. \(\large\bf{0=3(-2)+b}\)

bill533:

b = 6

bill533:

the answer is D ?

563blackghost:

You got the right intercept, but you changed the sign for some reason...it would be A.

bill533:

oh ok

bill533:

This graph represents the path of a ball rolling on a table. https://learning.k12.com/content/enforced/437405-COF_ID125380/G-GPE.B.5%20Q4.JPG?_&d2lSessionVal=x98CnccW19mQcrEo1ZJp3A15K A second ball rolls on the table parallel to the first ball. The path passes through (-2, 1). What is the equation of the path of the second ball? A. y = 1 B. x = 1 C. x = -2 D. y = -2

563blackghost:

What do you think your answer is?

bill533:

B

563blackghost:

Not quite, if `x=1` it would create a vertical slope, we want a slope that is parallel, so it would be `y=1`.

bill533:

Mario drew this line. https://learning.k12.com/content/enforced/437405-COF_ID125380/G-GPE.B.5%20Q3.JPG?_&d2lSessionVal=x98CnccW19mQcrEo1ZJp3A15K He wants to draw a second line that is parallel to the first and passes through (4, -5). What is the equation of Mario’s second line? A. y = −4/3x + 1/3 B. y = 3/4x − 8 C. y = −4/3x − 1/3 D. y = 3/4x + 8

563blackghost:

First identify the equation of the line. It passes through point `(0,7)` and has a rise of 3 with a run of 4 positively. So your equation is: \(\bf{y=\frac{3}{4}x+7}\) To find the parallel line you know that it has the same slope so: \(\bf{y=\frac{3}{4}x+b}\) Plug in your point to find b. \(\large\bf{-5=\frac{3}{4}(4)+b}\) What is b?

bill533:

b = -8

563blackghost:

Correct. So your equation is \(\large\bf{y=\frac{3}{4}x+8}\).

563blackghost:

welp -8

bill533:

that was incorrect

563blackghost:

\(\large\bf{y=\frac{3}{4}x-8}\) miss type im sorry

563blackghost:

nuuuuuuuuuuuuu

563blackghost:

i mistyped dat